If a discount of 20% off the retail price of a desk saves Mark $45, how much did he pay for the desk?


Found 3 solutions by nyc_function, Alan3354, w3suhl33:
Answer by nyc_function(2741)

0.20 * 45 = 9
Then 9 + 45 = $54

Answer by Alan3354(69443)

If a discount of 20% off the retail price of a desk saves Mark $45, how much did he pay for the desk?
---------------
The 20% = $45
He paid the remaining 80%, which is $180

Answer by w3suhl33(2)

1) Find retail price by re-writing as a equation with 20% in decimal form and x representing the retail price


.2x = 45
.2x/.2 = 45/.2
x = 225


2) Subtract discount from retail price


225-45 = 180


3) Answer:


Mark paid $180 for the desk after receiving a 20% discount from the retail price